OpenStudy (anonymous):

9/20 what is the eqivalent ratio

Directrix (directrix):

Other than 1, what are the common factors of 9 and 20?

OpenStudy (anonymous):

nothing else

Directrix (directrix):

Okay. So there are no equivalent ratios equal to 9/20 with numerator less than 9 and denominator less than 20. There are, however, infinitely many equivalent ratios equal to 9/20. 9/20 times 2/2 = 18/40 9/20 times 3/3 = 27/60 and so on. Check your problem instructions to see if you are looking for a simplifed equivalent ratio or for any equivalent ratio.